OpenStudy (anonymous):

6 cans of soda and 2 bags of chips cost $7.08 8 cans of soda and 5 bags of chips cost $12.45 What is the price per can of soda and bag of chips?

OpenStudy (saifoo.khan):

6x + 2y = 7.08 8x + 5y = 12.45

OpenStudy (saifoo.khan):

Solve simultaniously!

OpenStudy (saifoo.khan):

should i give u the answers?

OpenStudy (anonymous):

im trying.....lol

OpenStudy (saifoo.khan):

lol, Alright.

OpenStudy (anonymous):

ok, im confused

OpenStudy (chaise):

let s = soda let c = ships 6s + 2c = 7.08 8s + 5c = 12.45 30s + 10s = 35.40 16s + 10s = 24.90 Take away them from eachother 14s = 10.5 s = 0.75 6s+2c = 7.08 6x0.75+2c = 7.08 4.5 +2C = 7.08 2C = 2.58 c = 1.29

OpenStudy (anonymous):

would that be 14x + 7y = 19.53?

OpenStudy (saifoo.khan):

She got that right!

OpenStudy (saifoo.khan):

Little Angel, no!

OpenStudy (chaise):

If you need further explanation on anything I did in that equation I would be happy to explai.

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

where did you come up with the 30 + 10 equations?

OpenStudy (chaise):

I should say 30s+10c and 16s+10c. Sorry I mutliplied every term by 5, and then in the bottom question I multiplied every term by 2. Then as you can see the 10s are the same. I took the top equation away from the bottom equation and this left 10c = 0 I could then solve for c and got an answer

OpenStudy (saifoo.khan):

Chasieeee Owes a Good answer!

OpenStudy (anonymous):

I think I understand how you did it, thank you so so so much, I will be posting a few other problems as well, you are the first one to explain why you did something, thank you again